Calculate the discount factors (present value factors) at 8% for years 1, 2, 3, and 4 (show all work, including equations). Given your answer, aside for the math, why are the discount count factors declining? Must that be true? Explain fully.
Answer #1

Present Value Factor = 1/(1+r)t

r = interest rate

t = time period

PV factor for year 1 = 1/1.08 = 0.9259

PV factor for year 2 = 1/(1.08)2 = 0.8573

PV factor for year 3 = 1/(1.08)3 = 0.7938

PV factor for year 4 = 1/(1.08)4 = 0.7349

Discount factors are declining due to the concept of Time Value of money.

Value of a dollar today will decrease to a value of 0.9259 a year after, this is because the money available at present is worth more than money in a future period due to its earning potential which is 8% here.